H jet polarimeter status report
1 / 19

H-Jet polarimeter status report - PowerPoint PPT Presentation

  • Uploaded on

H-Jet polarimeter status report. Yousef Makdisi RSC 07/28/06. Jet operations in `06. Igor and Dima provided the new code to readout the full waveform and monitoring program. Ron provided the daily PC technical support and monitoring.

I am the owner, or an agent authorized to act on behalf of the owner, of the copyrighted work described.
Download Presentation

PowerPoint Slideshow about 'H-Jet polarimeter status report' - kasi

An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.

- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript
H jet polarimeter status report

H-Jet polarimeter status report

Yousef Makdisi

RSC 07/28/06

Jet operations in 06
Jet operations in `06

  • Igor and Dima provided the new code to readout the full waveform and monitoring program.

  • Ron provided the daily PC technical support and monitoring.

  • Anatoli and Yousef, the daily jet oversight and maintenance (a long haul but only 3 midnight calls)

  • T. Wise came to the rescue of the SFT unit.

  • MCR Operations took full responsibility and saw to it that data were collected in each fill.

Waveforms new h jet data format this year
Waveforms(new H-Jet data format this year)


Half maximum

Baseline = 8

Data collected @ 100 gev
Data Collected @ 100 GeV

Fills Dates Beam Events

7630-7652 3/18-25 B 2.5M

7662-7697 3/26-4/4 Y 2.8M

7718-7745 4/4-11 B 1.6M

4/11 Lost Si detector # 3

7780-7802 4/12-5/2 Y 3.8M

7810-7858 5/3-15 B 4.2M

7887-7946 5/19-6/1 Y 2.5M

6/2 Si detector acted up: reduced the bias from 200 – 175 V.

7949-8002 6/3-5 B 2.0M

Data collected at 31 2 gev
Data Collected at 31.2 GeV

Fills Dates Beam Events

8005-8054 6/9-17 B 3.6M

8055-8061 6/18-20 Y 2.6M

Data analysis
Data Analysis

  • The `05 data is being analyzed by Oleg.

  • Hiromi will complete the 24 GeV analysis for data taken in `04. This represents the first look at injection albeit with low statistics.

  • The 31.2 GeV data provides a good statistical sample taken in conjunction with the polarimeters. It is closest to calibration at injection. (No one??)

  • The 100 GeV `06 is waiting for a good volunteer.

  • The change in the DAQ to readout the full WFD waveform (Dima) does not allow the use of the old analysis programs.

  • Igor wrote the online analysis and monitoring program which could be used as a first step.

Operational issues
Operational issues

  • Nozzle blockage frequency: every 10-12 days. It takes 3 hours to warm up, ½ hour beam down, and two hours to cool down and back online.

  • Slower intensity ramp up than before. 3-4 days to reach full intensity, plateau for a few days and then a slow decrease to blockage.

  • Midway replaced the nozzle which improved matters somewhat.

  • The SFT phase became unstable for a period. Fixed by T. Wise by increasing the gain.

  • Lost some precious time due to memory full condition.

  • Lost one silicon detector, and another acted up.

Operations continued
Operations continued

  • Failed to take data with two beams at the same time:

  • Requirement that the two beams be separated by 4-6 mm.

  • The vertical collimation occluded the silicon acceptance.

  • Determined no polarization loss with 112 bunch operation and 1.1011 p/bunch implying the holding field uniformity is adequate.

  • Installed a CCD camera to look at light emitted as the beam hits the jet. This serves as another vertical beam emittance device. Our interest is to measure the molecular hydrogen contamination.

  • No pump failure this run.

What is next
What is next?

  • Determine the cause of the silicon failure.

  • Replace the failed detectors with existing spares.

  • Start discussion with Hamamatsu to purchase new detectors.

  • Purchase spare WFD components.

  • Will attempt to open up the vertical collimation to allow simultaneous two-beam measurement.

  • Wisconsin is building two RF cavities to allow a polarized deuteron jet beam.

  • Attempt to improve the jet vacuum to keep it on the beam line during Au-Au running.